When "0" is the exponent, how are the numbers effected if the exponent is inside or outside of the parenthesis. Example: (-1/4) with an exponent of zero inside and with an exponent of zero outside the parenthesis)?
Thanks!

Answer by stanbon(57307)

If inside (-1/4^0) = (-1/1) = -1
If outside (-1/4)^0 = 1
Cheers,
stan H.